OpenStudy (anonymous):

How to write .4 as a fraction

OpenStudy (dtan5457):

4/10

OpenStudy (dtan5457):

.4 is 40% also known as 4/10

OpenStudy (anonymous):

.4=4/10=2/5 @Droper

OpenStudy (michele_laino):

plase note that: \[0.4=\frac{ 0.4 }{ 1 }\] now, please multiply oth numerator anddenominator by 10, and yo will get your answer

OpenStudy (anonymous):

4/10, you can put the decimal as the numerator and 1 as the denominator, then you multiply by 10 on the numerator and the denominator. Then you just simplify! Hope this helps!!